OpenStudy (anonymous):

Partial Fraction decomposition: (x^2)/(x^4-2x^2-8)

OpenStudy (anonymous):

can you factor the denominator?

OpenStudy (anonymous):

Yes i factored denominator: (x^2+2)(x-2)(x+2)

OpenStudy (anonymous):

ok then we have \[\frac{x^2}{(x^2+2)(x-2)(x+2)}\] \[=\frac{Ax+B}{x^2+2}+\frac{C}{x-2}+\frac{D}{x-2}\] as a fist step

OpenStudy (anonymous):

actually i think in this case \(A=0\) but it doesn't need to be

OpenStudy (anonymous):

would (A/x-2) + (B/x+2) + ((Cx+D)/(x^2+2)) be correct as well?

OpenStudy (anonymous):

now add \[Ax+B(x+2)(x-2)+C(x^2+2)(x-2)+D(x^2+2)(x+2)=x^2\]

OpenStudy (anonymous):

yes of course

OpenStudy (anonymous):

it doesn't matter since addition is commutative but since i spent all that time writing it in latex, would you mind using it so i don't have to start over?

OpenStudy (anonymous):

Yes that's completely fine. thank you

OpenStudy (anonymous):

so now there are a couple of ways to proceed the easiest way is probably to replace \(x\) by \(2\) and see what you get

OpenStudy (anonymous):

the first two terms drop out, you are left with \[D(2^2+2)(2+2)=2^2\] and you can solve that for \(D\)

OpenStudy (anonymous):

that is really the basic gimmick is it clear what i did? you get \[D=\frac{1}{6}\] pretty much in your head

OpenStudy (anonymous):

one minute i am reading your work

OpenStudy (anonymous):

ok take your time i am going to get a snack if it is not clear, ask

OpenStudy (anonymous):

I understand what you did. I worked the problem and got D=1/6

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

now we can use the same gimmick only replace \(x\) yb \(-2\)

OpenStudy (anonymous):

Ok, so that cancels out the first and last term. I can now solve for C. C=-1/6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and now I plugin C and D into the first equation to solve for A & B?

OpenStudy (anonymous):

you can do that, but first lets think

OpenStudy (anonymous):

Ok..

OpenStudy (anonymous):

the first part is \[(Ax+B)(x+2)(x-2)\]

OpenStudy (anonymous):

if you multiply that out mentally you get an \(Ax^3\) term there is not \(x^3\) so \(A=0\)

OpenStudy (anonymous):

Sorry I do not understand why Ax^3 = 0

OpenStudy (anonymous):

because the numerator is \(x^2\)

OpenStudy (anonymous):

no \(x^3\) in it

OpenStudy (anonymous):

that means \(A=0\)

OpenStudy (anonymous):

So the exponent in the denominator cannot be greater than x^2?

OpenStudy (anonymous):

look at the original question

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

or even just this part \[(Ax+B)(x+2)(x-2)+C(x^2+2)(x-2)+D(x^2+2)(x+2)=x^2\]

OpenStudy (anonymous):

on the right hand side of the equal sign, there is only a little lonely \(x^2\) no \(x^3\) at all

OpenStudy (anonymous):

on the left side there will be, when you multiply out, one \(x^3\) term it will be \(Ax^3\)

OpenStudy (anonymous):

So A must equal 0 in order for the equation to be true

OpenStudy (anonymous):

ooh no scratch that, i am totally wrong, sorry

OpenStudy (anonymous):

No worries.

OpenStudy (anonymous):

there are other \(x^3\) terms there i forgot about them

OpenStudy (anonymous):

so yeah, substitute back what you know to solve for \(A\) and\(B\)

OpenStudy (anonymous):

you will find that \(A=0\) in any case

OpenStudy (anonymous):

\[x^2 = Ax+B(x+2)(x-2) - (1/6)(x^2+2)(x-2)+(1/6)(x^2+2)(x+2)\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yeah, now lets find \(B\)

OpenStudy (anonymous):

without doing a raft of solving a system of equations, which is the other long and tedious method of doing this

OpenStudy (anonymous):

\[(Ax+B)(x+2)(x-2)-\frac{1}{6}(x^2+2)(x-2)+\frac{1}{6}(x^2+2)(x+2)=x^2\]

OpenStudy (anonymous):

what will the constant be? the term without any \(x\) or any \(x^2\)?

OpenStudy (anonymous):

if it is not clear what i mean, let me know

OpenStudy (anonymous):

Ok i will work on the problem and let you know.

OpenStudy (anonymous):

Ok, I am confused. I am stuck on the part where you said "what will the constant be? the term without any x"

OpenStudy (anonymous):

from here \[-\frac{1}{6}(x^2+2)(x-2)\] when you mutiply out, you will get a number without an \(x\) namely \[-\frac{1}{6}\times 2\times (-2)=\frac{2}{3}\]

OpenStudy (anonymous):

from here \[\frac{1}{6}(x^2+2)(x+2)\] you will also get \(\frac{2}{3}\)

OpenStudy (anonymous):

Ok. That makes sense.

OpenStudy (anonymous):

from here \[(Ax+B)(x+2)(x-2)\] then number will be \(4B\)

OpenStudy (anonymous):

]scratch that it will be \[-4B\]

OpenStudy (anonymous):

Ok makes sense.

OpenStudy (anonymous):

on the right hand side there is no constant at all,

OpenStudy (anonymous):

that means \[\frac{2}{3}+\frac{2}{3}-4B=0\]or \[4B=\frac{4}{3}\] making \[B=\frac{1}{3}\]

OpenStudy (anonymous):

Ok. Why does the right side equal 0 though?

OpenStudy (anonymous):

ok i was mistaken before when i was trying to explain why \(A=0\) but now i am not the original numerator is just \(x^2\) it has no constant in it i

OpenStudy (anonymous):

the constant is 1

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Sorry I mean coefficient. yes there is no constant

OpenStudy (anonymous):

right

OpenStudy (anonymous):

you got \[(Ax+B)(x+2)(x-2)-\frac{1}{6}(x^2+2)(x-2)+\frac{1}{6}(x^2+2)(x+2)=x^2\] the contant on the left is \[\frac{4}{3}-4B\] and on the right it is \(0\)

OpenStudy (anonymous):

Makes sense now.

OpenStudy (anonymous):

that makes \(B=\frac{1}{3}\)

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

You made the constants equal to each other.

OpenStudy (anonymous):

So I get the answer: [(1/3)/(x^2+2)] + [(-1/6)/(x+2)] + [(1/6)/(x-2)] ?

OpenStudy (anonymous):

yes sometimes you can get away without doing that, but not in this case because of the \(x^2+2\) which is never zero so you can't use the previous trick of substitution

OpenStudy (anonymous):

that is one way to write it, yes

OpenStudy (anonymous):

Another way to write it would be [1/3(x^2+2)] - [1/6(x+2)] + [1/6(x-2)] ?

OpenStudy (anonymous):

So would my answer that I put be correct as well?

OpenStudy (anonymous):

normally you would write \[\frac{1}{3(x^2+2)}-\frac{1}{6(x+2)}+\frac{1}{6(x-2)}\]

OpenStudy (anonymous):

it is the same only in this topic do you see compound fractions allowed as an answer, but yes, both are the same

OpenStudy (anonymous):

Ok that makes sense. Thank you so much for you time I appreciate it. Good luck with your studies

OpenStudy (anonymous):

yw